14-05 In the following independent case, the company closes its books on December 31

1. Carla Co. sells $491,000 of 8% bonds on March 1, 2020. The bonds pay interest on September 1 and March 1. The due date of the bonds is September 1, 2023. The bonds yield 12%. Give entries through December 31, 2021.


Prepare a bond amortization schedule using the effective-interest method for discount and premium amortization. Amortize premium or discount on interest dates and at year-end. (Round answers to 0 decimal places, e.g. 38,548.)

Prepare all of the relevant journal entries from the time of sale until December 31, 2021. (Assume that no reversing entries were made.) (Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answers to 0 decimal places e.g. 58,971. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)
Answer #1

1. Semiannual coupon = $ 491,000 x 8 % x 1/2 = $ 19,640

i = 12 % x 1/2 = 6 %.

n = 7

PVA 6%, n=7 = [ { 1 - ( 1/1.06 )7 } / 0.06 ] = 5.5824

PV 6%, n= 7 = ( 1 / 1.06 ) 7 = 0.6651

Present value of the bonds = $ 19,640 x 5.5824 + $ 491,000 x 0.6651 = $ 109,638.34 + $ 326,564.10 = $ 436,202.44

Bond Amortization Table:

Date Amount Paid Interest Expense Discount Amortized Bond Carrying Value
3/1/20 436,202
9/1/20 19,640 26,172 6,532 442,734
3/1/21 19,640 26,564 6,924 449,658
9/1/21 19,640 26,979 7,339 456,997
3/1/22 19,640 27,420 7,780 464,777
9/1/22 19,640 27,887 8,247 473,024
3/1/23 19,640 28,381 8,741 481,765
9/1/23 19,640 28,875 9,235 491,000

In the books of Carla Co. :

Date Account Titles Debit Credit
$ $
3/1/20 Cash 436,202
Discount on Bonds Payable 54,798
Bonds Payable 491,000
9/1/20 Interest Expense 26,172
Discount on Bonds Payable 6,532
Cash 19,640
12/31/21 Interest Expense 17,709
Discount on Bonds Payable 4,616
Interest Payable 13,093
3/1/21 Interest Expense 8,855
Interest Payable 13,093
Discount on Bonds Payable 2,308
Cash 19,640
9/1/21 Interest Expense 26,979
Discount on Bonds Payable 7,339
Cash 19,640
12/31/21 Interest Expense 18,280
Discount on Bonds Payable 5,187
Interest Payable 13,093